A rental agent for vacation houses in southern Spain earns a flat amount of base pay per month plus a 7.5% commission on rentals sold during that month. One month he earns $5575 on rental sales of $45,000. What is his base monthly pay?

Answer by fractalier(2101)

Well his total salary is base plus commission, T = B + C.
We are given his total earnings as $5575, and we know how to figure commissions as commission rate times sales, so we have
5575 = B + (.075)(45000)
Now solve for B
5575 = B + 3375
B = 2200